Suppose country A has a total population of 1200. There are 200 children and 1000 adults. Of the 1000 adults, 850 are employed. Of the remaining 150 adults, 100 are not employed but are actively seeking work, and the remaining 50 are not employed and not actively seeking work.

a. calculate and interpret the labor force participation rate.

b. Calculate the unemployment rate and interpret it.

Suppose of the 100 individuals who are unemployed, 20 are in the middle of a job search, 50 have skills that are outmoded for the current economy, and the remaining 30 are unemployed due to an economic downturn. .

c. How many people are naturally unemployed? Explain.

d. What is the natural rate of unemployment? Interpret this rate.
Answer #1

a) Labor force participation rate is the number of working population actively participating in laborforce divided by total working population eligible to participate in the workforce. The active participants are 850 and total eligible are 100 then the labor force participation rate is 850/1000 = 0.85

b) Unemployment rate is the rate of unemployed people divided by labor force .i.e. 100 / 950 i.e. 0.105

c) Naturally unemployed are those which are frictional, structural and surplus unemployment. hence 70 people are naturally unemployed

d) natural rate of unemployment is naturally unemployed people divided by total unemployed = 70/ 100 = 0.7